# Excitation function of femtoscopic Lévy source parameters of pion pairs in EPOS4

### v1

@@ 3.1 Source shape
Figure 1: Lévy index $\alpha$ from EPOS4 calculations as a function of transverse mass $m_{T}$ at different energies. The bands show the systematical uncertainties.
Figure 2: Lévy index $\alpha$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ in three bins of $k_{T}$. The bands show the systematical uncertainties.
@@ 3.2 Source size
The extracted source radii exhibit the expected decrease with increasing transverse momentum and an overall growth with increasing collision energy, reflecting the expansion dynamics of the system. The energy dependence of $R_{o}$, $R_{s}$, and $R_{l}$ provides a systematic baseline for comparison with future experimental femtoscopy results.
[…]
@@
Figure 3: Lévy radii from EPOS4 calculations as a function of transverse momentum $m_{T}$ at $\sqrt{s_{NN}}=7.7-200\penalty 10000\ GeV$. The bands show the systematical uncertainties.
Figure 4: Energy dependence of Lévy radii from EPOS4 calculations at $k_{T}=0.375-0.425$ GeV/$c$. The bands show the systematical uncertainties.
@@
Ref. [16] has highlighted the sensitivity of the Lévy scale difference $R_{\mathrm{diff}}^{2}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ to potential critical-point signatures in the QCD phase diagram. Motivated by this, we investigated the energy dependence of both $R_{\mathrm{diff}}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, as shown in Figs. 6 and 6. The quantity $R_{\mathrm{diff}}$ decreases with increasing $m_{T}$, consistent with the behavior of the individual femtoscopic radii. And its collision-energy dependence shows a non-monotonic pattern, exhibiting a noticeable drop at 14.5 GeV, but without a clear global trend across the full energy range. The ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, on the other hand, exhibits a stonger overall decrease with increasing collision energy, accompanied by a notable pronounced enhancement in the interval 14.53914.5-39 GeV. The investigation of these observables for different equations of state is of crucial importance and presents the focus of subsequent studies.
@@
Figure 5: The radius difference $R_{\mathrm{diff}}^{2}$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ in three $k_{T}$ bins. The bands show the systematical uncertainties.
@@
Figure 6: The radius ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ at $k_{T}$=0.375-0.525GeV/cGeV/c. The bands show the systematical uncertainties.
@@ 3.3 Correlation strength
The transverse-mass dependence of the extracted $\lambda$ parameter at various collision energies is presented in Fig. 8, showing an overall increase with $m_{T}$. As shown in Fig. 8, $\lambda$ reveals a decreasing trend with increasing $\sqrt{s_{NN}}$, although noticeable anomalies are observed at $\sqrt{s_{NN}}=9.2$ and $11.5$ GeV, where the values deviate from the general behavior.
@@
Figure 7: Correlation strength parameters $\lambda$ from EPOS4 as a function of transverse momentum $m_{T}$ at $\sqrt{s_{NN}}=7.7-200\penalty 10000\ GeV$. The bands show the systematical uncertainties.
Figure 8: Energy dependence of $\lambda$ from EPOS4 at $k_{T}=0.375-0.525\penalty 10000\ GeV/c$. The bands show the systematical uncertainties.
3.4 Comparison to data and EPOS3
@@
In this analysis, the femtoscopic parameters of pions from $200$ GeV Au+Au collisions are generally consistent with those obtained from EPOS3 in Ref. [25] within approximately $2\sigma$, with the notable exception of the $R_{\mathrm{side}}$ radius, which exhibits a systematic reduction. For $R_{\mathrm{side}}$, EPOS4 predicts systematically smaller values, with deviations corresponding to $N_{\sigma}=2.5$$5.8$, where $N_{\sigma}={\Delta}/{\sigma_{\mathrm{combined}}}$ and $\Delta$ is the difference between the EPOS4 and EPOS3 results, while $\sigma_{\mathrm{combined}}=\sqrt{\sigma_{\mathrm{EPOS4}}^{2}+\sigma_{\mathrm{EPOS3}}^{2}}$ denotes the quadrature sum of their uncertainties. Due to this similarity to EPOS3, and the detailed data comparisions of EPOS3 to data of Ref. [7] in Ref. [25], it can be stated that EPOS4 is also compatible with the data.
Due to its particular importance [7], we nevertheless discuss the data comparision in case of the $\lambda$ parameter. As mentioned, Ref. [25] provided a detailed discussion showing that the experimental intercept parameter $\lambda$ can be regarded as the integral of $D_{\mathrm{cc}}$, which represents the core-core component of the particle-emitting source, when the total number of pairs is normalized to unity. It is important to note however, that pairs at large distances consist of pions produced far away from the interaction point, thus their “distance of closest approach” would be sizeable—resulting in its removal from the actual pion sample in an experimental analysis. Thus in fact it is the number of pion pairs up to a given distance that constitute the “full” pion sample. In order to make a direct comparison with the experimental results, similarly to Ref. [25], different integration limits $\rho_{\lambda}^{max}$ (0.1 mm, 1 cm, and 50 cm) were applied in this work to obtain $\lambda$, which was then compared with the PHENIX experimental data at $\sqrt{s_{NN}}=200$ GeV [7]. The $m_{\mathrm{T}}$ dependences of $\lambda$ and the scaled $\lambda/\lambda_{\rm max}$ are shown in Fig. 10 and Fig. 10, respectively. While the nominal $\lambda$ value can be described only if choosing a small $\rho_{\lambda}^{max}$, the description of the scaled correlation strength measurements by EPOS4 is adequate in all cases.
Figure 9: Correlation strength parameters $\lambda$ from EPOS4 for variotus choices of $\rho_{\lambda}^{max}$ (lines of different colors with bands), compared to PHENIX data of Ref. [7] (black points), as a function of mT. The bands shows the systematical uncertainties of the EPOS4 calculations. The statistical and systematical uncertainties of the PHENIX data points are shown as error bars and boxes, respectively.
Figure 10: Normalized correlation strength parameters $\lambda$$\lambda_{max}$ from EPOS4 for various choices of $\rho_{\lambda}^{max}$ (lines of different colors with bands), compared to PHENIX data of Ref. [7] (black points), as a function of mT. The bands show the systematical uncertainties of the EPOS4 calculations. The statistical and systematical uncertainties of the PHENIX data points are shown as error bars and boxes, respectively.
@@ 4 Summary and outlook
We presented detailed investigations of the simulated particle emitting source in $\sqrt{s_{NN}}=7.7-200$ GeV 0-10% centrality Au+Au collisions within the framework of the EPOS4 model. We calculated the spatial pair distribution $D(\vec{\rho})$ for identical charged pions in the longitudinally comoving system, for several bins of pair transverse mass ($m_{T}$) and calculated their one-dimensional projections in the Bertsch-Pratt out, side, and long coordinates. We fitted these projections simultaneously with projections of a single 3D Lévy distribution, and found a statistically and qualitatively also adequate description of the source, up to distances of a few dozen (or up to a hundred) femtometers. We then extracted the following source parameters: Lévy index (or shape parameter) $\alpha$, spatial Lévy scales (or HBT radii) $R_{\rm out,side,long}$, and correlation strength parameter $\lambda$. We then investigated the $m_{T}$ and $\sqrt{s_{NN}}$ dependence of these source parameters.
We found that the radii $R_{\rm out,side,long}$ decrease with transverse mass at all collision energies, a characteristic observation attributed to the expansion of the system. This decrease appeared to be the strongest in the longitudinal direction. The radii show a moderate increase with collision energy, also strongest for the longitudinal direction. An average scale ($R_{\rm avg}$) and out-side anisotropy (in form of the difference and the ratio of the out and side radii) were also investigated. The anisotropy appears to decrease towards higher collision energies. One of the important future research directions is to investigate this for a modified equation of state, as the change in the order of the phase transition is expected to have characteristic signature in the increase of the correlation radius in the out direction ($R_{\rm out}$).
The Lévy exponent, describing the source shape, was also investigated as a function of transverse mass and collision energy. A weak but smooth increase with transverse mass is apparent for all collision energies. A less smooth decrease is found as a function of collision energy. As the $\alpha$ parameter is expected to be strongly affected by the equation of state, in particular a decrease in $\alpha$ is expected at the critical point, again it is one of the crucial research directions to calculate $\alpha$ versus $\sqrt{s_{NN}}$ for an equation of state describing a second order phase transition.
@@
The correlation strength parameter $\lambda$, extracted based on the integral of the fitted Lévy distribution, and corresponding to the core fraction among all pion pairs, was also investigated as a function of transverse mass and collision energy. A clear increase is found with transverse mass at low $m_{T}$, for all collision energies. This can also be described as a “hole” at small $m_{T}$, as discussed by PHENIX in Ref. [7]. EPOS4, similarly to EPOS3 [25], describes the scaled correlation strength $\lambda/\lambda_{\rm max}$ PHENIX data well. Furthermore, a weak decrease is found in $\lambda$ for larger collision energies.
It is important to investigate the equation of state dependence of these observables, and compare them to data, once Lévy source parameter measurements at several collision energies will be available.
5 Acknowlegements
@@
This research was funded by the NKFIH grants TKP2021-NKTA-64, PD-146589, K-146913, K-138136, and NKKP-152097.

### v2

@@ 3.1 Source shape
Figure 2: Lévy index $\alpha$ from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of transverse mass $m_{T}$ at different energies. The bands show the systematic uncertainties.
Figure 3: Lévy index $\alpha$ from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of collision energy $\sqrt{s_{NN}}$ in three bins of $k_{T}$. The bands show the systematic uncertainties. The vertical blue dashed line at 19 GeV denotes the low-energy region where EPOS4 is expected to be less reliable [18].
@@ 3.2 Source size
The extracted source radii exhibit the expected decrease with increasing average transverse momentum of the pair and an overall growth with increasing collision energy, reflecting the expansion dynamics of the system. The energy dependence of $R_{\mathrm{out}}$, $R_{\mathrm{side}}$, and $R_{\mathrm{long}}$ provides a systematic baseline for comparison with future experimental femtoscopy results.
As shown in Fig. 4, the femtoscopic radii $R_{\mathrm{out}}$, $R_{\mathrm{side}}$, and $R_{\mathrm{long}}$ exhibit a clear decrease with increasing $m_{T}$. In addition, Fig. 5 presents their energy dependence for the $k_{T}$ bin of $0.375-0.425~\rm{GeV}/c$} (corresponding to $m_{T}\approx 0.422~\rm{GeV}/c^{2}$}), including the averaged radius $R_{\mathrm{avg}}=\sqrt{(R_{\mathrm{out}}^{2}+R_{\mathrm{side}}^{2}+R_{\mathrm{long}}^{2})/3}$}. It can be seen that $R_{\mathrm{long}}$ and $R_{\mathrm{side}}$ increase with collision energy. The increase of $R_{\mathrm{long}}$ is consistent with a longer system lifetime at higher collision energies. In EPOS4, this originates from the higher initial energy density and the correspondingly longer hydrodynamic evolution. In contrast, $R_{\mathrm{out}}$ exhibits a rather weak dependence on energy. This behavior was already observed in the STAR data  [33]. The gradual increase of $R_{\mathrm{long}}$ and $R_{\mathrm{side}}$ with collision energy, reflecting spatial expansion at higher energies, is further illustrated by the pair distance distributions shown in Fig. 1. We show $D(\rho_{\mathrm{out}}^{\mathrm{LCMS}})$, $D(\rho_{\mathrm{side}}^{\mathrm{LCMS}})$, and $D(\rho_{\mathrm{long}}^{\mathrm{LCMS}})$ for the $k_{T}$ bin $0.225-0.275~\rm{GeV}/c$, corresponding to $m_{T}\approx 0.285~\rm{GeV}/c^{2}$}), at low (7.7 GeV), intermediate (19.6 GeV), and high (200 GeV) energies. A clear widening of $\rho_{\mathrm{long}}$ and $\rho_{\mathrm{side}}$ distributions is observed, indicating an increase of the longitudinal and transverse source sizes with energy, while the $\rho_{\mathrm{out}}$ distributions show no significant change, consistent with the weak energy dependence of $R_{out}$ discussed above.
@@
Figure 4: Lévy radii from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of transverse mass $m_{T}$ at $\sqrt{s_{NN}}=7.7-200~\rm{GeV}$}. The bands show the systematic uncertainties.
Figure 5: Energy dependence of Lévy radii from EPOS4 simulations of $0-10\%$ Au+Au collisions at $k_{T}=0.375-0.425~\rm{GeV}/c$}. The bands show the systematic uncertainties. The vertical blue dashed line at 19 GeV denotes the low-energy region where EPOS4 is expected to be less reliable [18].
@@
Ref. [16] has highlighted the sensitivity of the Lévy scale difference $R_{\mathrm{diff}}^{2}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ to potential critical-point signatures in the QCD phase diagram. Motivated by this, we investigated the energy dependence of both $R_{\mathrm{diff}}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, as shown in Fig. 6 and Fig. 7. The quantity $R_{\mathrm{diff}}$ decreases with increasing $m_{T}$, consistent with the behavior of the individual femtoscopic radii. And its collision-energy dependence shows a non-monotonic pattern, exhibiting a noticeable drop at 14.5 GeV followed by partial recovery and a gradual decrease, yielding no clear global trend across the full energy range. The ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, on the other hand, exhibits a stronger overall decrease with increasing collision energy, accompanied by a notable pronounced enhancement in the interval 14.53914.5-39 GeV. The strong deviation of this ratio from unity, especially at the lowest energies, is probably stemming from a not fully adequate description of several details, such as pre-thermalized acceleration, equation of state, or viscous corrections [34]. The investigation of these observables for different equations of state is of crucial importance and presents the focus of subsequent studies.
@@
Figure 6: The radius difference $R_{\mathrm{diff}}^{2}$ from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of collision energy $\sqrt{s_{NN}}$ in three $k_{T}$ bins. The bands show the systematic uncertainties. The vertical blue dashed line at 19 GeV denotes the low-energy region where EPOS4 is expected to be less reliable [18].
@@
Figure 7: The radius ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of collision energy $\sqrt{s_{NN}}$ at $k_{T}=0.375-0.525~\rm{GeV}/c$}. The bands show the systematic uncertainties. The vertical blue dashed line at 19 GeV denotes the low-energy region where EPOS4 is expected to be less reliable [18].
@@ 3.3 Correlation strength
The transverse mass dependence of the extracted $\lambda$ parameter at various collision energies is presented in Fig. 8, showing an overall increase with $m_{T}$. As shown in Fig. 9, $\lambda$ reveals a decreasing trend with increasing collision energy, although noticeable anomalies are observed at $\sqrt{s_{NN}}=9.2$ and $11.5$ GeV, where the values deviate from the general behavior.
@@
Figure 8: Correlation strength parameters $\lambda$ from EPOS4 simulations of $0-10\%$ Au+Au collisions as a function of transverse mass $m_{T}$ at $\sqrt{s_{NN}}=7.7-200~\rm{GeV}$. The bands show the systematic uncertainties.
Figure 9: Energy dependence of $\lambda$ from EPOS4 simulations of $0-10\%$ Au+Au collisions at $k_{T}=0.375-0.525~\rm{GeV}/c$. The bands show the systematic uncertainties. The vertical blue dashed line at 19 GeV denotes the low-energy region where EPOS4 is expected to be less reliable [18].
3.4 Comparison with EPOS3 and experimental data
@@
In this analysis, the femtoscopic parameters of pions from $200$ GeV Au+Au collisions are generally consistent with those obtained from EPOS3 in Ref. [25] within approximately $2\sigma$, with the notable exception of the $R_{\mathrm{side}}$ radius, which exhibits a systematic reduction. For $R_{\mathrm{side}}$, EPOS4 predicts systematically smaller values, with deviations corresponding to $N_{\sigma}=2.5-5.8$}, where $N_{\sigma}={\Delta}/{\sigma_{\mathrm{combined}}}$ and $\Delta$ is the difference between the EPOS4 and EPOS3 results, while $\sigma_{\mathrm{combined}}=\sqrt{\sigma_{\mathrm{EPOS4}}^{2}+\sigma_{\mathrm{EPOS3}}^{2}}$} denotes the quadrature sum of their uncertainties. The $N_{\sigma}$ deviations between EPOS4 and EPOS3 for each Lévy parameter, evaluated in different $m_{T}$ bins, are listed in Table 2. Due to this similarity to EPOS3, and the detailed data comparison of EPOS3 to data of the one-dimensional measurement [7] in Ref. [25], it can be stated that EPOS4 is also compatible with the data. However, it will be important to compare our predictions with three-dimensional measurements, once these are performed.
Table 2: $N_{\sigma}$ significance of the differences between EPOS4 and EPOS3 (EPOS4 - EPOS3) for the Lévy parameters at different transverse mass in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV.
[…]
0.24 1.8 -0.095 0.11 -5.8 -1.3 -2.3
0.28 1.3 -0.12 -0.37 -5.3 -0.96 -1.6
0.33 2.1 -0.084 -0.25 -4.7 -0.74 -1.3
0.38 2.3 -0.044 0.34 -3.8 -0.66 -1.1
0.42 1.8 -0.088 1.2 -3.6 -0.39 -0.75
0.47 1.0 -0.088 1.5 -3.1 -0.31 -0.64
0.52 1.5 -0.15 1.8 -2.8 -0.25 -0.57
0.57 0.20 -0.056 1.7 -2.8 -0.068 -0.49
0.62 0.70 -0.061 2.0 -2.6 -0.23 -0.54
0.66 0.12 0.031 2.8 -2.5 -0.26 -0.52
0.71 0.68 0.025 3.6 -2.5 -0.36 -0.58
@@ 4 Summary and outlook
We present a detailed investigations of the simulated particle emitting source in 0-10% central Au+Au collisions at collision energies $\sqrt{s_{NN}}=7.7-200$ GeV within the framework of the EPOS4 model. The spatial pair distance distribution $D(\vec{\rho})$ for identical charged pions was calculated in the longitudinally comoving system for several bins of pair transverse mass, and its one-dimensional projections in the Bertsch-Pratt out, side, and long coordinates were obtained. These projections were simultaneously fitted with those of a single 3D Lévy distribution, providing a statistically and qualitatively adequate description of the source up to distances of a few dozen (or up to a hundred) femtometers. The following source parameters were extracted: Lévy index (or shape parameter) $\alpha$, spatial Lévy scales (or HBT radii) $R_{\rm out,side,long}$, and correlation strength parameter $\lambda$. And then the transverse mass $m_{T}$ and energy dependence of these source parameters was investigated.
@@
We found that the radii $R_{\rm out,side,long}$ decrease with $m_{T}$ at all collision energies, a characteristic observation attributed to the expansion of the system. This decrease appeared to be the strongest in the longitudinal direction. The radii show a moderate increase with collision energy, also strongest for the longitudinal direction. An average scale ($R_{\rm avg}$) and out-side anisotropy (in form of the difference and the ratio of the out and side radii were) also investigated. The anisotropy appears to decrease towards higher collision energies. The Lévy exponent, describing the source shape, was also investigated as a function of $m_{T}$ and collision energy. A weak but smooth increase with $m_{T}$ is observed across all energies, whereas a less regular decrease occurs with increasing collision energy. The correlation strength parameter $\lambda$, extracted based on the integral of the fitted Lévy distribution, and corresponding to the core fraction among all pion pairs, was also investigated as a function of transverse mass and collision energy. A clear increase is found with $m_{T}$ for all collision energies. This can also be described as a “hole” at small $m_{T}$, as discussed by PHENIX in Ref. [7]. In EPOS4, the predicted $\lambda$ values at $\sqrt{s_{NN}}$ = 200 GeV are very close to those obtained in EPOS3 [25], which were shown to describe the PHENIX data well. Furthermore, a weak decrease is found in $\lambda$ for larger collision energies.
One important future research direction is to investigate the equation of state dependence of the Lévy source observables. In particular, a modified equation of state, where the order of the phase transition changes, is expected to leave characteristic signatures in the increase of the correlation radius in the out direction ($R_{\rm out}$). Similarly, the $\alpha$ parameter is expected to be modified due to criticality, with a decrease near the critical point [32], making it essential to calculate $\alpha$ versus collision energy for an equation of state describing a second-order phase transition, as well as for the case of enhanced non-hydrodynamic modes. Finally, we note that it will be crucial to investigate the equation of state dependence of these observables, and compare them to data, once Lévy source parameter measurements at several collision energies will be available.
Acknowledgements
@@
This research was funded by the NKFIH grants TKP2021-NKTA-64, PD-146589, K-146913, K-138136, and NKKP ADVANCED 152097.